OpenStudy (anonymous):

Need help on geting started please... 2x^2/3-5x^1/3-3=0

jimthompson5910 (jim_thompson5910):

let z = x^(1/3) square both sides to get z^2 = [ x^(1/3) ]^2 z^2 = x^(2/3)

jimthompson5910 (jim_thompson5910):

so the original equation turns into 2z^2 - 5z - 3 = 0 use the quadratic formula to solve for z to get z = ??? or z = ???